### v1

# Learning Value of Information towards Joint Communication and Control in 6G V2X

Abstract
As Cellular Vehicle-to-Everything (C-V2X) evolves towards future sixth-generation (6G) networks, Connected Autonomous Vehicles (CAVs) are emerging to become a key application. Leveraging data-driven Machine Learning (ML), especially Deep Reinforcement Learning (DRL), is expected to significantly enhance CAV decision-making in both vehicle control and V2X communication under uncertainty. These two decision-making processes are closely intertwined, with the value of information (VoI) acting as a crucial bridge between them. In this paper, we introduce Sequential Stochastic Decision Process (SSDP) models to define and assess VoI, demonstrating their application in optimizing communication systems for CAVs. Specifically, we formally define the SSDP model and demonstrate that the MDP model is a special case of it. The SSDP model offers a key advantage by explicitly representing the set of information that can enhance decision-making when available. Furthermore, as current research on VoI remains fragmented, we propose a systematic VoI modeling framework grounded in the MDP, Reinforcement Learning (RL) and Optimal Control theories. We define different categories of VoI and discuss their corresponding estimation methods. Finally, we present a structured approach to leverage the various VoI metrics for optimizing the “When”, “What”, and “How” to communicate problems. For this purpose, SSDP models are formulated with VoI-associated reward functions derived from VoI-based optimization objectives. While we use a simple vehicle-following control problem to illustrate the proposed methodology, it holds significant potential to facilitate the joint optimization of stochastic, sequential control and communication decisions in a wide range of networked control systems.
Index Terms:
Vehicle-to-everything (V2X), 6G, autonomous driving, reinforcement learning (RL), and Value of Information (VoI).
I Introduction
Connected Autonomous Vehicles (CAVs) integrate automated control with advanced communication technologies to achieve self-driving capabilities with minimal or no human intervention. Cellular Vehicle-to-Everything (C-V2X) technologies are advancing rapidly and are on the brink of commercial deployment. As C-V2X evolves towards six-generation (6G), CAVs are anticipated to be a key application in the 6G era [1].
Unlike Autonomous Vehicles (AVs) without communication capabilities, CAVs can achieve a more accurate and comprehensive perception of the driving environment by leveraging critical information exchanged through Vehicle-to-Everything (V2X) communications, which is referred to as ‘V2X information’ in this paper. This enables them to make more reliable and efficient driving decisions by integrating V2X information with local sensory data [2].
Recent years have witnessed a growing interest in applying data-driven Machine Learning (ML) to CAVs. Deep Reinforcement Learning (DRL), which combines deep learning (DL) with reinforcement learning (RL), has shown significant advantages over traditional model-driven control and optimization methods, as it requires fewer restrictive assumptions about the stochastic properties of system dynamics [3, 4, 5, 6]. Leveraging DRL for both vehicle control and V2X communication, an intelligent, integrated decision-making system presents a promising solution for ensuring safe and efficient autonomous driving (AD). To enable this, Markov Decision Process (MDP) models must be conceived for framing the vehicle control and V2X communication problems, respectively. The system would be comprised of two types of agents - a control agent and a communication agent - that employ DRL algorithms to solve the MDP models and learn optimal policies for decision-making. Notably, the control and communication decisions may operate at different time scales [7].
The control agent makes decisions on vehicle control inputs based on available sensory and V2X information. These inputs are then executed by actuators, such as steering, throttle, brakes, and gear shifting, enabling autonomous vehicle operation. It is widely recognized that providing the control agent with more information at the time of decision-making can lead to better decisions. However, an overwhelming amount of information can be toxic to the control agent and greatly increases the computational complexity of making decisions due to the curse of dimensionality problem [8]. Ideally, the control agent would have access to the necessary and sufficient information to make optimal decisions. In practice, however, information comes at a cost and can be imperfect. This paper mainly focuses on the communication costs incurred for information dissemination and the imperfections in information resulting from non-ideal communication.
Given the challenges of perfectly delivering all available V2X information to CAVs, a series of key questions arise: what information should be transmitted, when should it be transmitted, and how should the limited radio resources be scheduled for transmission? Thus, the communication agent is responsible for making these decisions.
Understanding the importance of sharing specific information with other CAVs is crucial for making effective communication decisions. To evaluate this, the value of information (VoI) is studied in this paper, as it quantifies the expected benefits of acquiring additional information for decision-making [9]. Although VoI has been investigated across diverse research domains, each with its unique perspective [10], this paper specifically focuses on VoI in the context of integrated control and communication.
Since the information is utilized by control agents for decision-making, the VoI should be evaluated by these agents. The communication agents - whether at the base station (BS), roadside infrastructure, CAVs and other information sources - can leverage the estimated VoI, alongside traditional communications performance metrics such as throughput and delay, to optimize their communication decisions.
The VoI depends on how the control agent reacts to missing or imperfect information by adjusting its decisions, which in turn relies on the type of controllers used. Traditionally, AV controllers have been designed based on classical control theory, such as linear controller, $\mathcal{H}_{\infty}$ controller, and Sliding Mode Controller (SMC) [11]. In this paper, we focus on VoI when DRL-based AV controllers are employed. Our results, however, extend to any AV controller designed by solving MDPs, including both model-based dynamic programming/optimal control and model-free reinforcement learning methods.
The MDP model describes a dynamic system whose ‘state’ evolves over time under the influence of ‘actions’ taken at discrete instances of time. The outcome of each action is captured by a scalar reward. The objective is to derive a policy function that maps states to actions, enabling an agent to make a sequence of decisions (on actions) over time that maximizes the expected cumulative rewards.
In MDP formalism, the ‘state’ is a set of variables that contain essential information for decision-making. However, there is a lack of thorough discussion regarding the formal definition of state in the existing literature. Specifically, what information should it incorporate? Considering the communication costs and limitations for information dissemination, the above question can be expanded to the following three questions:
-Q1
What information can enhance decision-making if made available?
-Q2
How can the benefit of information be measured? In other words, how can VoI be defined and estimated?
-Q3
How can VoI be leveraged to make communication decisions for information dissemination and radio resource management (RRM)?
I-A What information can enhance decision-making if made available?
I-A1 Related Works
The only requirement for a state in an MDP model is that it satisfies the Markov property. In other words, given the current state and action, the reward and the next state are independent of past states and actions. Silver points out that there are actually two versions of state, namely the environment state and the agent state [12]. The environment state is the environment’s private state representation, which is usually not fully visible to the agent. It includes the information the environment uses to determine the next state and reward, while some of this information may be irrelevant to decision-making. Meanwhile, the agent’s policy is a function of the agent state, which consists of the information the agent uses to select the action at the time of decision-making. This state is the agent’s internal representation of the environment state. Since the information in the environment state may not be fully and accurately accessible to the agent, the agent state is constructed based on the information available to it. If the agent state does not satisfy the Markov property, the resulting decision process becomes a partial observable MDP (POMDP), which can be reformulated into an MDP by incorporating the history of past states and actions into the agent state.
Although Silver’s definition distinguishes between the environment state and the agent state, it remains unclear which information in the environment state is irrelevant to decision-making. More importantly, it is not well-defined what information could enhance decision-making if incorporated into the agent state. In contrast, Powell provides a formal definition of the state [13]: “A state variable is the minimally dimensioned function of history that is necessary and sufficient to compute the decision (policy) function, the transition function, and the contribution (reward) function.” He further explains that the definition provides a quick test for the validity of a state variable: “If there is a piece of data in either the policy function, the transition function, or the reward function that is not in the state variable, then we do not have a complete state variable.” However, although the transition and reward functions are objective models of the environment, the policy function is subjective and defined by the agent, with the state as its input argument. This leads to a potential deadlock in using Powell’s definition to identify a state: we need to determine the policy function to define the state, yet we also need to ascertain the state to establish the policy function.
I-A2 Contribution 1
We introduce a general model known as the Sequential Stochastic Decision Process (SSDP) model, of which the MDP model is a special case. Based on the SSDP model, we provide a mathematically rigorous and practical method to identify the set of information that can enhance decision-making if made available. Specifically, we define exogenous information within the SSDP model as any information that affects the transition function, the reward, or both, but is unavailable prior to decision-making. We demonstrate that this exogenous information can improve decision-making once it becomes available. Moreover, if the exogenous information is not available, any information that can predict the exogenous information can also enhance decision-making. Any other information apart from the above is not useful. We also clarify the relationship between our findings and the definitions of state provided by Silver and Powell.
I-B How can the benefit of information be measured? In other words, how can VoI be defined and estimated?
I-B1 Related Works
From the discussion above, we understand that the agent state may contain only partial information about the environment state. While the agent can utilize its historical data to construct a Markov state, this enhanced state may still be insufficient to fully represent the environment state. In this case, how do we measure the benefit of acquiring additional information that exists in the environment state but is not included in the agent state? Moreover, some information in the agent state may be imperfect due to factors such as transmission delay when delivered to the agent. Thus, how do we measure the benefit of knowing perfect information instead of imperfect information?
Information theory provides powerful tools for quantifying the amount of uncertainty in the value of a random variable, and has had a profound influence on the design of communication systems. However, as Howard indicates in his seminal work, “no theory that involves just the probabilities of outcomes without considering their consequences could possibly be adequate in describing the importance of uncertainty to a decision maker” [9]. Howard introduces the theory of VoI, and asserts that numerical values can be assigned to the elimination or reduction of any uncertainty. To explain the concept and intuition of VoI, Howard uses a bidding problem as an example, where the expected profit of eliminating any certainty is used as the VoI. However, Howard’ work does not provide a rigorous definition or mathematical formula for calculating VoI. Following his work, various definitions of VoI have been proposed for different applications and objectives. A survey on VoI in Wireless Sensor Network (WSN) and Internet of Things (IoT) by Alawad et al. [10] provides a structured literature review on the fragmented works in this area. Despite these contributions, a theoretical foundation for VoI is still lacking, which would allow VoI to be estimated using uniform formulas for different applications.
I-B2 Contribution 2
We provide a systematic method for defining and estimating VoI in the context of making a sequence of decisions under uncertainty. This approach has broad practical applications, with autonomous driving serving as a key example. Our definition of the VoI is rooted in MDP and RL theories, and is derived from the SSDP model we develop.
We categorize VoI into two types: (1) utility-based VoI, which measures the performance loss in terms of expected cumulative rewards in SSDP models; and (2) information theory-based VoI, which quantifies the VoI based on how much better we can predict the next state and reward. The rationale for defining information theory-based VoI stems from our assertion in Contribution 1 that only information influencing the state transition and reward contributes to enhanced decision-making.
For utility-based VoI, we extend Howard’s VoI concept and define two versions: Value of Missing Information (VoMI) and Value of Imperfect Information (VoPI). Additionally, we differentiate between Expected Cumulative VoI (EVoI) and Immediate VoI (IVoI) to address the cases when a sequence of decisions are made based on the missing/imperfect information versus when only a single decision is made based on this information. The mathematical relationship between EVoI and IVoI is also established.
To derive the information theory-based VoI, we calculate the conditional KL divergence between the probability distribution given by the product of the transition models of the original state and the V2X information, and the probability distribution given by the transition model of the augmented state, which inclues both the original state and the V2X information. While information theory-based VoI is easier to derive given the model of the environment, utility-based VoI has several important advantages. Therefore, we introduce a DRL-based method to estimate the utility-based VoI by offline training policies with missing/imperfect and full/perfect information, respectively. Then, the regret and advantage function can be estimated to quantify the EVoI and IVoI, respectively.
I-C How to leverage the VoI to make communication decisions for information dissemination and RRM?
I-C1 Related Works
VoI has been used to guide communication decisions in existing studies [14, 15, 16, 17]. The general idea is to transmit only information with high VoI or prioritizing the transmission of such information. However, VoI in these studies is typically calculated using intuitive metrics, such as the importance of information and its timeliness. Therefore, the vague definition of VoI may lead to misinformed communication decisions.
In V2X communication, there is a growing interest aimed at designing RRM mechanisms to minimize the Age of Information (AoI) [18, 19, 20, 21], rather than focusing solely on traditional metrics such as throughput and delay. While AoI measures the importance of information by assessing its timeliness, it does not account for the broader benefits that information can provide. In contrast, VoI directly quantifies the advantage of information in enhancing the performance of specific tasks. Unfortunately, there is currently a lack of studies focused on VoI-aware RRM.
In MDP and RL communities, the problems of “What to Communicate” and “When to Communicate” have been studied in the context of Decentralized-POMDP (Dec-POMDP) [22, 23] and Multi-agent RL (MARL) [24, 25]. The goal in these studies is to simultaneously learn both a communication policy and an acting policy to maximize performance while adhering to communication constraints. However, VoI is rarely used in this line of research [23]. One negative consequence of this neglect is that it is challenging to explicitly express the trade-off between maximizing the performance of considered task and that of traditional communication services. As a result, communication constraints in these studies are often simplified, typically involving restrictions on the number of messages or the agents allowed to transmit.
I-C2 Contribution 3
We propose a systematic framework to leverage VoI for optimizing communication systems. We classify communication decisions into three broad categories and identify the three time scales for decision-making. A general optimization objective is formulated to make VoI-oriented communication decisions, aiming to maximize communication performance for traditional services while minimizing performance degradation in vehicle control due to non-ideal communications. This performance degradation is quantified by the EVoI.
In our framework, SSDP models are developed for the “When”, “What”, and “How” to Communicate problems, with IVoI-associated reward functions derived from the optimization objective. Additionally, we detail the application of various VoI metrics in guiding different types of communication decisions, providing a structured approach to optimize system-level performance.
The reminder of this paper is organized as follows. Section II describes the system model. Section III provides a brief introduction to the preliminaries of MDP. Section IV presents the SSDP model as a theoretical foundation, addressing Q1. Subsequently, Sections V and VI address Q2 by introducing the VoI definition and the DRL-based estimation method, respectively. Section VII addresses Q3 by presenting a VoI-oriented optimization framework for communication systems. Finally, Section VIII concludes the paper with a discussion on future research directions.
II System Model
In this paper, we consider discrete time system where the time is discretized into equal-length control intervals indexed by $k\in\mathcal{K}=\{0,1,\cdots,\infty\}$. Without loss of generality, we focus on the ego CAV shown in Fig. 1. The ego CAV is equipped with a transmitter and receiver pair, multiple local on-board sensors (such as cameras, LiDARs and radars) and actuators, as well as a pair of control and communication agents. There are numerous remote sensors deployed on other vehicles and within the infrastructure. These remote sensors can transmit their collected data to the ego CAV using V2X communication, either directly via Vehicle-to-Vehicle (V2V) links or through a Base Station (BS) using Vehicle-to-Infrastructure (V2I) links. A communication agent operates at the BS, and potentially at some of the remote sensors.
Figure 1: Illustration of system model
At each control interval $k$, the control agent determines the CAV control input $u_{k}$ [^1] We use $x_{k}$ to represent any variable $x$ at the beginning of control interval $k$. based on the available sensory and V2X information. The control input $u_{k}$ is applied to the actuators at the beginning of control interval $k$ and holds constant within the time period of the control interval.
Due to communication limitations in practical systems, the V2X information received by the control agent may be missing or imperfect. For illustration purposes, Fig. 2 presents examples of typical scenarios, which are explained as follows.
Figure 2: Examples of typical scenarios
Scenario A: Missing Information
At control interval $k$, a piece of V2X information is not received by the ego CAV. The missing information can be caused by
  1. 1.
    the remote sensors withholding the information without transmission to save the usage of radio resources, or
  2. 2.
    transmission loss in the communication systems.
In this case, the ego CAV could either
  • replace the missing information with some dummy bits; or
  • replace the missing information with the last received information.
The latter case converts the missing information to delayed information as described below.
Scenario B: Imperfect Information
The V2X information received by the ego CAV could be imperfect. Two main categories of imperfect information are listed below.
  1. 1.
    Delayed Information: At control interval $k$, the ego CAV receives delayed information due to transmission latency in the communication systems, or in cases where information is missing, it is replaced with the last received data, as previously described. This means that the information used by the control agent to make decisions at control interval $k$ was sampled at control interval $k-\tau_{k}$ by the remote sensors. $\tau_{k}$ is normally a random variable whose distribution depends on the stochastic properties of the wireless channel and RRM mechanisms.
  2. 2.
    Imprecise Information: At control interval $k$, a noisy version of the information is received by the CAV. Specifically, the received information may deviate from the transmitted information due to
    • the remote sensors processing the information before transmission and thus results in information distortion, or
    • transmission errors in the communication systems.
Given the vast amount of remote sensory information and the impracticality of sharing all of it perfectly via V2X communication, this paper explores the answers to the three questions identified in Section I.
Case 1.
We consider a simple vehicle-following control problem with two vehicles, wherein the position, velocity and acceleration of a following vehicle (follower) $i$ are denoted by $p_{i}$, $v_{i}$, $acc_{i}$, respectively. Here $p_{i}$ represents the one-dimensional position of the center of the front bumper of vehicle $i$.
The vehicle $i$ obeys the dynamics model described by a first-order system:
$\dot{p}_{i}=v_{i},$
$\dot{v}_{i}=acc_{i},$
$\dot{acc}_{i}=-\frac{1}{\rho_{i}}acc_{i}+\frac{1}{\rho_{i}}u_{i},$ (1)
where $\rho_{i}$ is a time constant representing the driveline dynamics of vehicle $i$.
We denote the headway of vehicle $i$ at time $t$, i.e., bumper-to-bumper distance between $i$ and its preceding vehicle (predecessor) $i-1$, by $d_{i}$, which satisfies
$d_{i}=p_{i-1}-p_{i}-L_{i-1},$ (2)
where $L_{i-1}$ is the the length of vehicle $i-1$.
According to Constant Time-headway Policy (CTHP), vehicle $i$ aims for maintaining a desired headway of $d_{\sigma,i}$, given by
$d_{c,i}=c_{i}+h_{i}v_{i},$ (3)
where $c_{i}$ is a constant standstill distance for vehicle $i$ and $h_{i}$ is the desired time-gap of vehicle $i$.
The control errors, including position error $e_{pi}$ and velocity error $e_{vi}$ are defined as
$e_{pi}=d_{i}-d_{r,i},$
$e_{vi}=v_{i-1}-v_{i}.$ (4)
Let $x_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k}]^{\mathrm{T}}$, where the subscript $k$ indicates the values of the variables at the beginning of control interval $k$. The system dynamics in discrete time are derived from (1) on the basis of forward Euler discretization:
$x_{i,k+1}=A_{i}x_{i,k}+B_{i}u_{i,k}+C_{i}acc_{i-1,k},$ (5)
where
$A_{i}=\begin{bmatrix}1&T&-h_{i}T\\ 0&1&-T\\ 0&0&1-\frac{T}{\tau_{i}}\end{bmatrix},B_{i}=\begin{bmatrix}0\\ 0\\ \frac{T}{\tau_{i}}\end{bmatrix},C_{i}=\begin{bmatrix}0\\ T\\ 0\end{bmatrix}.$ (6)
At the beginning of each control interval, the control agent of vehicle $i$ has to determine the control input $u_{i,k}$ based on the observation of the state. The velocity $v_{i,k}$ and acceleration $acc_{i,k}$ can be measured locally, while the control errors $e_{pi,k}$ and $e_{vi,k}$ can be quantified by a radar unit mounted at the front of the vehicle. On the other hand, vehicle $i$ can only obtain the acceleration $acc_{i-1,k}$ of the predecessor through V2X communications. To determine the optimal action $u_{i,k}$ for $k=0,1,\cdots,K-1$, we must decide what, when, and how to share information by V2X communications from predecessor $i-1$ to its follower $i$.
III Preliminaries - Markov Decision Process
Since our focus is on AV controllers developed through MDP solutions, we provide an overview of MDP fundamentals to facilitate understanding of the subsequent contents in this paper.
An MDP is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $S_{k}\in\mathcal{S}$ and $a_{k}\in\mathcal{A}$ are state and action [^2] The CAV control input $u_{k}$ is the action $a_{k}$ in the AV control problem. at time step $k$ within state space $\mathcal{S}$ and action space $\mathcal{A}$, respectively. At each time step $k$, the action $a_{k}$ is determined by the agent based on the state $S_{k}$ and applied to the environment. Partly as a result of the $a_{k}$, a numerical reward signal $R_{k+1}\in\mathcal{R}$ is generated and the state $S_{k}$ transits to a new state $S_{k+1}\in\mathcal{S}$ at time step $k+1$. Both the next state and the reward are related to the current state and action, where a well-defined discrete probability distribution can be given as $\mathrm{Pr}(S_{k+1},R_{k+1}|S_{k},a_{k})$ when the set of states, actions, and rewards all have a finite number of elements. The four-argument probability distribution defines the dynamics of the MDP, based on which the other functions of the environment can be derived, such as the state transition probability
$\mathrm{Pr}(S_{k+1}|S_{k},a_{k})=\sum_{r\in\mathcal{R}}\mathrm{Pr}(S_{k+1},R_{k+1}=r|S_{k},a_{k}),$ (7)
and the two-argument reward function
$r(S_{k},a_{k})$ $=\mathrm{E}[R_{k+1}|S_{k},a_{k}]$
$=\sum_{r\in\mathcal{R}}r\sum_{s^{\prime}\in\mathcal{S}}\mathrm{Pr}(S_{k+1}=s^{\prime},R_{k+1}=r|S_{k},a_{k}).$ (8)
The state transition matrix $P$ defines state transition probabilities from all states $S_{k}\in\mathcal{S}$ to all successor states $S_{k+1}\in\mathcal{S}$.
The agent uses a policy $\pi$ to determine the action $a_{k}$, where the policy is a conditional probability distribution over the action space given a state
$\pi(a|s)=\mathrm{Pr}(a_{k}=a|S_{k}=s),\forall s\in\mathcal{S},a\in\mathcal{A}$ (9)
The goal of the agent is to learn a policy that maximizes the cumulative reward it receives in the long run. There are different settings for formulating the cumulative reward in the theories of dynamic programming and RL, such as episodic setting, discounted setting, and average-reward setting. In this paper, we use a unified notation that covers both the episodic and discounted settings for explanatory purposes [26], while our core ideas can be extended to other settings as well. To formally define the optimization objective of the agent, the state-value or value function $v_{\pi}(s)$ is defined as the expected cumulative reward from starting in state $s$ and then following policy $\pi$
$v_{\pi}(s)=\mathrm{E}_{\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}R_{k+l+1}\arrowvert S_{k}=s\right],\forall s\in\mathcal{S},$ (10)
where $\gamma\in[0,1]$ is the discount factor, and $\mathrm{E}_{\pi}[\cdot]$ denotes the expected value of a random variable when the agent follows policy $\pi$.
Similarly, the action-value or Q function $q_{\pi}(s,a)$ is defined as the expected cumulative reward from starting in state $s$, taking action $a$, and then following policy $\pi$
$q_{\pi}(s,a)=\mathrm{E}_{\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}R_{k+l+1}\arrowvert S_{k}=s,a_{k}=a\right],\forall s\in\mathcal{S},a\in\mathcal{A}$ (11)
Therefore, the objective of the control agent can be defined as deriving the optimal policy $\pi^{*}$ that maximizes the value function for all the states
$v_{\pi^{*}}(s)=\sup_{\pi}v_{\pi}(s),\forall s\in\mathcal{S}.$ (12)
To quantify how good a policy is, the performance $J_{\pi}$ of a policy $\pi$ is defined as
$J_{\pi}=\mathrm{E}_{p(S_{0})}[v_{\pi}(S_{0})]=\mathrm{E}_{\pi,p(S_{0})}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}\arrowvert S_{0}\right],$ (13)
where $p(S_{0})$ denotes the probability distribution of the initial state $S_{0}$.
Based on the value function $v_{\pi}(s)$ and Q function $q_{\pi}(s,a)$, we can derive the advantage function as
$A_{\pi}(s,a)=q_{\pi}(s,a)-v_{\pi}(s),$ (14)
which quantifies whether action $a$ is better or worse than the policy’s default behavior.
For AV control tasks, state $S_{k}$ can be constructed based on the available sensory information. However, some useful information for decision-making may be missing from or imperfect in the state. Therefore, we aim to quantify the benefit of making decisions when a piece of missing or imperfect information becomes available or perfect in the state, and we seek answers to Questions 1 and 2 outlined in Section I.
IV Sequential Stochastic Decision Process
The MDP model faces some limitations in addressing Q1 and Q2 in Section I. One key issue is that the MDP model does not explicitly represent the useful information missing from the state. To develop a rigorous and practical method for defining and evaluating the VoI, we first extend the MDP model to a more general model known as the SSDP model, based on which the Q1 is addressed.
IV-A SSDP Basics
Definition 1 (Sequential Stochastic Decision Process).
Define an SSDP by a tuple $\mathfrak{S}=\{\mathcal{S},\mathcal{A},\mathcal{W},f^{S},r,f^{W},\gamma\}$, where $\mathcal{S}$, $\mathcal{A}$, and $\gamma$ are the state space, action space, and discount factor as defined for an MDP, while
  • $W_{k}\in\mathcal{W}$ represents the exogenous information within its outcome space $\mathcal{W}$ that is not available at time step $k$ before the decision $a_{k}$ is made. The exogenous information $W_{k}$ includes any information that meets at lease one of the following two conditions: (1) $W_{k}$ affects the transition of state $S_{k}$, and (2) $W_{k}$ affects the reward $R_{k+1}$;
  • $f^{S}$ is the system’s state transition function governing $S_{k+1}=f^{S}(S_{k},a_{k},W_{k})$;
  • $R_{k+1}=r(S_{k},a_{k},W_{k})$ is the reward function;
  • $f^{W}$ is the function governing the exogenous information $W_{k}=f^{W}(\tilde{W}_{k})$, where $\tilde{W}_{k}\in\tilde{\mathcal{W}}$ represents all the deterministic and/or random parameters that affect the value of $W_{k}$.
Case 2.
In Case 1, the vehicle-following control problem operating without V2X communications can be formulated as an SSDP $\{S_{i,k},a_{i,k},W_{i,k},f^{S_{i}},f^{W_{i}},R\}$ with
  • state $S_{i,k}=x_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k}]^{\mathrm{T}}$;
  • action $a_{i,k}=u_{i,k}$;
  • exogenous information $W_{i,k}=acc_{(i-1),k}$;
  • state transition function $f^{S_{i}}$ governing
    $S_{i,k+1}=f^{S_{i}}(S_{i,k},a_{i,k},W_{i,k}),$ (15)
    which is given by (5) and (6);
  • exogenous information function $f^{W_{i}}$ given by
    $W_{i,k}$ $=f^{W_{i}}\left(W_{i,(k-1)},u_{(i-1),(k-1)}\right)$
    $=(1-\frac{1}{\rho_{i-1}})W_{i,(k-1)}+\frac{1}{\rho_{i-1}}u_{(i-1),(k-1)},$ (16)
    which is derived from the predecessor $i-1$’s dynamic model $\dot{acc}_{i-1}=-\frac{1}{\tau_{i-1}}acc_{i-1}+\frac{1}{\tau_{i-1}}u_{i-1}$ based on forward Euler discretization;
  • and the reward function $r(S_{i,k},a_{i,k})$ given by
    $r(S_{i,k},a_{i,k})=-\{(e_{pi,k})^{2}+\alpha(e_{vi,k})^{2}+\beta(a_{i,k})^{2}\},$ (17)
    where $\alpha$ and $\beta$ are the weights that are positive and can be adjusted to determine the relative importance of minimizing the position error, velocity error and the control input.
The SSDP is an extended version of MDP, incorporating the cases with non-Markov states. The following theorem provides the condition when the SSDP reduces to an MDP. The proof of Theorem 1 is provided in Appendix A.
Theorem 1 (Relationship between SSDP and MDP).
The SSDP reduces to an MDP if and only if $\tilde{W}_{k}\subseteq\{S_{k},a_{k},\xi_{k}\}$ in Definition 1, where $\xi_{k}$ is a set of random variables whose distribution does not depend on the past states, actions and exogenous information, i.e., $\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}$.
Case 3.
For the SSDP in Case 2, we have $\tilde{W}_{i,k}=\{W_{i,(k-1)},u_{(i-1),(k-1)}\}$. Therefore, this SSDP is not an MDP according to Theorem 1, since $\tilde{W}_{i,k}$ depends on the past exogenous information $W_{i,k-1}$.
IV-B Augmented-State SSDP
Given the SSDP in Definition 1, we aim to identify the set of additional information that can improve decisions beyond what is already included in the state. We start by inspecting the exogenous information $W_{k}$ in Definition 1, and examine whether its availability before decision making leads to better decisions.
Definition 2 (Augmented-state SSDP with $W_{k}$).
Assume that the exogenous information $W_{k}$ in the SSDP given in Definition 1 is available before decision $a_{k}$ is made. Then we define an augmented-state SSDP by $\tilde{\mathfrak{S}}=\{\tilde{\mathcal{S}},\mathcal{A},\tilde{\mathcal{W}},f^{\tilde{S}},r,f^{\tilde{W}},\gamma\}$, where the augmented state $\tilde{S}_{k}=(S_{k},W_{k})$ is obtained by including the exogenous information $W_{k}$ in the enlarged state at time step $k$. The action $a_{k}$, the reward function $r(\tilde{S}_{k},a_{k})=r(S_{k},a_{k},W_{k})$, and the discount factor $\gamma$ are the same as those of the original SSDP.
The system’s state transition function $f^{\tilde{S}}$ becomes
$\tilde{S}_{k+1}$ $=\begin{pmatrix}S_{k+1}\\ W_{k+1}\end{pmatrix}$
$=\begin{pmatrix}f^{S}(S_{k},a_{k},W_{k})\\ f^{W}(\tilde{W}_{k+1})\end{pmatrix}$
$=f^{\tilde{S}}(\tilde{S}_{k},a_{k},\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}).$ (18)
From (2), the exogenous information becomes $\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}$, where $\tilde{W}_{k+1}\in\tilde{\mathcal{W}}$. Finally, the exogenous information function becomes $\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}=f^{\tilde{W}}(\tilde{\tilde{W}}_{k+1})$, where $\tilde{\tilde{W}}_{k+1}\in\tilde{\tilde{\mathcal{W}}}$ represents all the parameters that affect the value of $\tilde{W}_{k+1}\backslash{\{\tilde{S}_{k},a_{k}\}}$.
Case 4.
The vehicle-following control problem relying on V2X communications where $acc_{(i-1),k}$ is transmitted from the preceding vehicle $i-1$ can be formulated as an augmented-state SSDP $\{\tilde{S}_{i,k},a_{i,k},\tilde{W}_{i,k},f^{\tilde{S}_{i}},f^{\tilde{W}_{i}},R\}$ with
  • state $\tilde{S}_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k},acc_{(i-1),k}]^{\mathrm{T}}=[(S_{i,k})^{\mathrm{T}},W_{i,k}]^{\mathrm{T}}$;
  • exogenous information $\tilde{W}_{i,k}=u_{(i-1),k}$;
  • state transition function $f^{\tilde{S}_{i}}$ given by
    $\tilde{S}_{i,k+1}$ $=\begin{pmatrix}S_{i,k+1}\\ W_{i,k+1}\end{pmatrix}=\begin{pmatrix}f^{S_{i}}(S_{i,k},a_{i,k},W_{i,k})\\ f^{W_{i}}(W_{i,k},u_{(i-1),k}),\end{pmatrix}$
    $=f^{\tilde{S}_{i}}(\tilde{S}_{i,k},a_{i,k},u_{(i-1),k})$ (19)
    where $f^{S_{i}}$ is formulated in (15), and $f^{W_{i}}$ is given in (2);
  • exogenous information function $f^{\tilde{W}_{i}}$ given by
    $u_{(i-1),k}=f^{\tilde{W}_{i}}(\tilde{\tilde{W}}_{i,k}).$ (20)
    where $\tilde{\tilde{W}}_{i,k}$ is the set of information that affects the control input $u_{(i-1),k}$ of predecessor $i-1$, such as the road conditions and the controller parameters of predecessor $i-1$.
Theorem 2.
The optimal policy of the augmented-state SSDP $\tilde{\pi}^{*}(\tilde{S}_{k})$ is at least as good as that of the original SSDP $\pi^{*}(S_{k})$.
The proof of Theorem 2 is provided in Appendix B. Theorem 2 demonstrates that including exogenous information $W_{k}$ in the augmented state $\tilde{S}_{k}$ has the potential to improve the decisions. Alternatively, in cases where $W_{k}$ is unavailable, the knowledge of $\tilde{W}_{k}$ can be leveraged to conceive an augmented-state SSDP, since $W_{k}=f^{W}(\tilde{W}_{k})$.
Definition 3 (Augmented-state SSDP with $\tilde{W}_{k}$).
Assume that $\tilde{W}_{k}$ instead of $W_{k}$ in the SSDP given in Definition 1 is available before decision $a_{k}$ is made. Then we define an augmented-state SSDP by $\tilde{\mathfrak{S}}^{\prime}=\{\tilde{\mathcal{S}}^{\prime},\mathcal{A},\tilde{\tilde{\mathcal{W}}},f^{\tilde{S}^{\prime}},r,f^{\tilde{\tilde{W}}},\gamma\}$, where the augmented state $\tilde{S}^{\prime}_{k}=(S_{k},\tilde{W}_{k})$ is obtained by including $\tilde{W}_{k}$ in the enlarged state at time step $k$. The action $a_{k}$, the reward function $r(\tilde{S}_{k},a_{k})=r(S_{k},a_{k},f^{W}(\tilde{W}_{k}))$, and the discount factor $\gamma$ are the same as those of the original SSDP.
The system’s state transition function $f^{\tilde{S}}$ becomes
$\tilde{S}_{k+1}$ $=\begin{pmatrix}S_{k+1}\\ \tilde{W}_{k+1}\end{pmatrix}$
$=\begin{pmatrix}f^{S}(S_{k},a_{k},f^{W}(\tilde{W}_{k}))\\ f^{\tilde{W}}(\tilde{\tilde{W}}_{k+1},S_{k},a_{k},f^{W}(\tilde{W}_{k}))\end{pmatrix}$
$=f^{\tilde{S}^{\prime}}(\tilde{S}^{\prime}_{k},a_{k},\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}).$ (21)
From (3), the exogenous information becomes $\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}$, where $\tilde{\tilde{W}}_{k+1}\in\tilde{\tilde{\mathcal{W}}}$. Finally, the exogenous information function becomes $\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}=f^{\tilde{\tilde{W}}}(\tilde{\tilde{\tilde{W}}}_{k+1})$, where $\tilde{\tilde{\tilde{W}}}_{k+1}\in\tilde{\tilde{\tilde{\mathcal{W}}}}$ represents all the parameters that affect the value of $\tilde{\tilde{W}}_{k+1}\backslash\{\tilde{S}^{\prime}_{k},a_{k}\}$.
The following Lemma 1 identifies the set of additional information that can improve decisions.
Lemma 1.
Given the SSDP in Definition 1, the set of information $\mathcal{I}_{k}\in\{W_{k},\tilde{W}_{k},\tilde{\tilde{W}}_{k},\cdots\}$ is beneficial for enhanced decision-making if becomes available before decision $a_{k}$ is made. Here, $W_{k}$ represents the exogenous information, $\tilde{W}_{k}$ represents the predictive information of $W_{k}$, and $\tilde{\tilde{W}}_{k}$ is the predictive information of $\tilde{W}_{k}$, and so on. Specifically, the availability of:
  • Exogenous information $W_{k}$,
  • Predicative information $\tilde{W}_{k}$ when $W_{k}$ is unavailable,
  • Predicative information $\tilde{\tilde{W}}_{k}$ when both $W_{k}$ and $\tilde{W}_{k}$ are unavailable, and so on
can improve decisions. Any information $I_{k}$ cannot enhance decision-making if $I_{k}\notin\mathcal{I}_{k}$, for all $\mathcal{I}_{k}\in\{W_{k},\tilde{W}_{k},\tilde{\tilde{W}}_{k},\cdots\}$.
Remark 1 (how to conceive the state $S_{k}$ in Definition 1).
$S_{k}$ is constructed based on the information available to the agent. However, not all information is necessarily relevant to decision-making. Lemma 1 provides a method to formulate the state in any MDP or SSDP model:
  • Step 1: Pick any available information that is necessary and sufficient to compute the reward $R_{k+1}$ and form $S_{k}$.
  • Step 2: For each piece of available information $I_{k}$:
    • if $I_{k}$ is necessary and sufficient to compute the next state $S_{k+1}$, update $S_{k}\leftarrow\{S_{k},I_{k}\}$.
Remark 2 (Relationship between Lemma 1 and the environment state).
The environment state introduced by Silver [12] is essentially the augmented state $\tilde{S}_{k}=(S_{k},W_{k})$ as defined in Definition 2, which incorporates the state $S_{k}$ and exogenous information $W_{k}$. Once the augmented state $\tilde{S}_{k}$ is known, the next state $S_{k+1}$ and reward $R_{k+1}$ are completely determined. All the information in the environment state is useful in making decisions, although some information may be more important than others. We will quantify the VoI in the environment state in Section IV. Given the environment state, no additional information can enhance decision-making. Therefore, the environment state includes the necessary and sufficient information for making the optimal decisions.
Remark 3 (Relationship between Lemma 1 and the state definition by Powell).
According to Lemma 1, Powell’s state definition can be revised to “A state variable is the minimally dimensioned function of history that is necessary and sufficient to compute the transition function and the reward function.” The decision or policy function is removed from the definition. Actually, if a piece of information is necessary and sufficient to compute the transition function and the reward function, it is useful to compute the policy function.
Notice that the augmented-state SSDPs in Definition 2 and Definition 3 consider that the full set of information $W_{k}$ or $\tilde{W}_{k}$ is available before decision making. In general, we can conceive an augmented-state SSDP to incorporate any information $I_{k}\in\mathcal{I}_{k}$ in the augmented state $\tilde{S}_{k}$.
IV-C Augmented-State SSDP with imperfect information
Given the SSDP in Definition 1, we consider a more complex scenario where a piece of information $I_{k}\in\mathcal{I}_{k}$ is available before decision making but could be imperfect. Let $I_{k}^{\rm ip}$ denote the imperfect information. As discussed in Section II, the imperfect information can be divided into (1) delayed information, i.e., $I_{k}^{\rm ip}=I_{k-\tau_{k}}$; and (2) imprecise information, i.e., $I_{k}^{\rm ip}=I_{k}+Z_{k}$, where $Z_{k}$ is a random disturbance.
In either case, we can conceive augmented-state SSDPs to leverage the imperfect information for making improved decisions. In the following, as an example, we formulate an augmented-state SSDP where the exogenous information $W_{k}$ is delayed.
Definition 4 (Random Delay SSDP).
Assume that the delayed exogenous information $W_{k-\tau_{k}}$ in the SSDP given in Definition 1 is available before decision $a_{k}$ is made. Then we define a random delay SSDP by $\tilde{\mathfrak{S}}^{\rm d}=\{\tilde{\mathcal{S}}^{\rm d},\mathcal{A},\tilde{\mathcal{W}}^{\rm d},f^{\tilde{S}^{\rm d}},r,f^{\tilde{W}^{\rm d}},\gamma\}$, where the augmented state $\tilde{S}_{k}^{\rm d}=(S_{k-\tau_{k}},W_{k-\tau_{k}},\{a_{k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}}^{k-1},\tau_{k})$ is obtained by including delayed exogenous information $W_{k-\tau_{k}}$, the last $\tau_{\mathrm{max}}$ actions $\{a_{k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}}^{k-1}$, as well as the observation delay $\tau_{k}$ in the enlarged state at time step $k$. The action $a_{k}$, the reward function $r(S_{k},a_{k},W_{k})$, and the discount factor $\gamma$ are the same as those of the original SSDP.
The system’s state transition function $f^{\tilde{S}}$ becomes
$\tilde{S}_{k+1}^{\rm d}$ $=\begin{pmatrix}\tilde{S}_{k+1-\tau_{k+1}}\\ \{a_{k^{\prime}}\}_{k^{\prime}=k+1-\tau_{\mathrm{max}}}^{k}\\ \tau_{k+1}\end{pmatrix}$
$=\begin{pmatrix}f^{S}(\cdots f^{S}(f^{S}(\tilde{S}_{k-\tau_{k}},a_{k-\tau_{k}},\tilde{W}_{k-\tau_{k}}),a_{k-\tau_{k}+1},\tilde{W}_{k-\tau_{k}+1})\cdots,a_{k-\tau_{k+1}},\tilde{W}_{k-\tau_{k+1}})\\ \{a_{k^{\prime}}\}_{k^{\prime}=k+1-\tau_{\mathrm{max}}}^{k-1},a_{k}\\ f^{\tau}(\tau_{k},W^{\tau}_{k})\end{pmatrix}$
$=f^{\tilde{S}}(\tilde{S}_{k}^{\rm d},a_{k},\tilde{W}_{k}^{\rm d}).$ (25)
From (4), the exogenous information becomes $\tilde{W}_{k}^{\rm d}=(\{\tilde{W}_{k^{\prime}}\}_{k^{\prime}=k-\tau_{k}}^{k-\tau_{k+1}},W^{\tau}_{k})\in\tilde{\mathcal{W}}^{\rm d}$. Finally, the exogenous information function is $\tilde{W}_{k}^{\rm d}=f^{\tilde{W}^{\rm d}}(\tilde{\tilde{W}}_{k}^{\rm d})$, where $\tilde{\tilde{W}}_{k}^{\rm d}\in\tilde{\tilde{\mathcal{W}}}^{\rm d}$ represents all the parameters that affect the value of $\tilde{W}_{k}^{\rm d}$.
Case 5.
A Random Delay SSDP can be formulated for with state
$\tilde{S}_{i,k}^{\rm d}=[x_{i,k-\tau_{i,k}},acc_{(i-1),k-\tau_{i,k}},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}}^{k-1},\tau_{i,k}]^{\mathrm{T}}$.
In the rest of the paper, we use $\tilde{S}^{\rm ip}$ to denote an augmented state with imperfect information. $\tilde{S}^{\rm d}_{k}$ defined in Definition 4 is an example of $\tilde{S}^{\rm ip}$.
V VoI Definition
In Section IV, we have identified the additional set of information beyond the state variables that empowers the control agent to enhance decision-making. In this section, we formally define the VoI to quantify the merit of information. Broadly speaking, two categories of VoI are introduced: utility-based VoI and information theory-based VoI.
V-A Utility-based VoI
V-A1 Value of Missing Information vs. Value of Imperfect Information
Since the ultimate objective for a control agent in obtaining information is to make improved decisions for achieving better control performance, the utility-based VoI measures the performance difference when decisions are made based on partial/imperfect versus full/perfect information.
Recall the state $S_{k}$ includes the set of information that is currently available to the agent. Consider the case when an additional piece of information $I_{k}\in\mathcal{I}_{k}$ becomes available before decision-making. Two scenarios are considered: (1) $I_{k}$ is perfectly available; (2) imperfect information $I_{k}^{\rm ip}$ is available instead of $I_{k}$.
We define two types of VoI, namely the Value of Missing Information (VoMI) and the Value of Imperfect Information (VoII), to measure the value of $I_{k}$ in these two scenarios, respectively.
Definition 5.
(Value of Missing Information (VoMI)) The VoMI is defined as the difference in the performance of the optimal policy $\pi^{*}$ for the SSDP, where the state $S_{k}$ does not include information $I_{k}$, and the performance of the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes $I_{k}$.
Definition 6.
(Value of Imperfect Information (VoII)) The VoII is defined as the difference in the performance of the optimal policy $\tilde{\pi}^{\rm ip*}$ for the augmented-state SSDP with imperfect information, where the state $\tilde{S}_{k}^{\mathrm{ip}}$ is constructed based on the imperfect information $I_{k}^{\rm ip}$, and the performance of the optimal policy $\tilde{\pi}^{*}$ for the augmented-state SSDP $\tilde{S}_{k}=\{S_{k},I_{k}\}$, where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes the perfect information $I_{k}$.
Note that both VoMI and VoII correspond to the performance difference between a policy with inferior performance and a policy with superior performance. The former policy is the optimal policy based on partial/imperfect observation (state with partial/imperfect information), while the latter policy is the optimal policy based on full/perfect observation (state with full/perfect information). Depending on how the performance difference is measured, both VoMI and VoII can be further divided into two categories: Expected Cumulative VoI and Immediate VoI, which are discussed in the next subsection.
V-A2 Expected Cumulative VoI vs. Immediate VoI
For ease of understanding, we use unified notations $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}$ to represent the inferior policy and superior policy in the VoMI and VoII definitions, respectively. Additionally, $S_{k}^{\rm{inf}}$ and $S_{k}^{\rm{sup}}$ are used to represent the partial/imperfect observation and full/perfect observation, respectively.
Definition 7.
(Expected Cumulative VoI (EVoI)) The EVoI $\bar{\Xi}$ is defined as the difference in the performance of the optimal policy $\pi^{\rm{inf}}$ based on the partial/imperfect observation and the optimal policy $\pi^{\rm{sup}}$ based on the full/perfect observation, respectively. The performance difference is quantified as
$\bar{\Xi}=J_{\pi^{\rm{inf}}}-J_{\pi^{\rm{sup}}},$ (22)
where both $J_{\pi^{\rm{inf}}}$ and $J_{\pi^{\rm{sup}}}$ are calculated according to (13).
The EVoI indicates the performance loss as a consequence of the control agent consistently receiving the partial/imperfect observations at every control interval. In the terminology of RL, EVoI corresponds to the regret of inferior policy $\pi^{\rm{inf}}$. It is worth noting that the EVoI is a non-positive value, with a larger EVoI indicating less performance loss. When the EVoI reaches its maximum value of zero, no performance loss is experienced.
Definition 8.
(Immediate Value of Information (IVoI)) Given the full/perfect observation $S_{k}^{\rm{sup}}$ and partial/imperfect observation $S_{k}^{\rm{inf}}$ at control interval $k$, the IVoI $\xi_{k}$ is defined as the difference in the performance starting from control interval $k$, where the decision at control interval $k$ is made by the optimal policy $\pi^{\rm{inf}}$ based on the partial/imperfect observation and the optimal policy $\pi^{\rm{sup}}$ based on the full/perfect observation, respectively, while the rest of the sequential control decisions are made by $\pi^{\rm{sup}}$ based on the full/perfect observations $S_{k+1}^{\rm{sup}},S_{k+2}^{\rm{sup}},\dots$, i.e.,
$\xi_{k}=A_{\pi^{\rm{sup}}}(S_{k}^{\rm{sup}},a_{k}^{\rm{inf}})|_{a_{k}^{\rm{inf}}\sim\pi^{\rm{inf}}(\cdotp|S_{k}^{\rm{inf}})},$ (23)
where $a_{k}^{\rm{inf}}\sim\pi^{\rm{inf}}(\cdotp|S_{k}^{\rm{inf}})$ indicates that the action $a_{k}^{\rm{inf}}$ is sampled according to $\pi^{\rm{inf}}$.
The IVoI indicates the performance loss as a consequence of the control agent receiving the partial/imperfect observations only at a specific control interval $k$, while consistently obtaining full/perfect observations throughout the remaining time. Moreover, IVoI is calculated for a specific full/perfect and partial/imperfect observation pair at control interval $k$. In other words, IVoI reveals that if the full/perfect observation at control interval $k$ is $S_{k}^{\rm{sup}}=s$, what is the performance loss if the agent only has access to the partial/imperfect observation $S_{k}^{\rm{inf}}$ instead? In the terminology of RL, IVoI corresponds to the advantage function of the optimal policy $\pi^{\rm{sup}}$. Similar to EVoI, IVoI is also a non-positive value, where maximizing IVoI leads to minimizing the performance loss.
It is proved in [27] that the EVoI is equivalent to the expected cumulative IVoI, similar to the relationship between the expected cumulative reward and immediate reward in RL.
Lemma 2.
The EVoI equals to the expected cumulative IVoI, i.e.,
[…]
where $\pi^{\rm CM}$ denotes the communication policy. Note that $\pi^{\rm CM}$ affects the stochastic properties of communication nonideality, such as delay distribution, which in turn impacts the transition of the partial/impaired observations received by the control agent. It is essential to clarify that the term Immediate VoI doesn’t imply its sole focus on quantifying the disparity in the immediate rewards at control interval $k$, when the control decisions are made based on the partial/imperfect versus full/perfect observations, respectively. Since the decision at control interval $k$ not only affects the immediate reward at the current control interval but also the future rewards through its impact on the next state, the IVoI measures the performance loss in terms of the expected cumulative reward just as the EVoI does.
Therefore, utility-based VoI can be categorized into two main types: EVoI and IVoI. Each of these types is further divided into two subcategories: VoMI and VoII. As illustrated in Table. I, this classification results in four distinct categories of utility-based VoI: Expected VoMI (EVoMI), Expected VoII (EVoII), Immediate VoMI (IVoMI), and Immediate VoII (IVoII). According to Definition 5 and 6, the policy/state/action $\pi^{\rm{inf}}$/$S_{k}^{\rm{inf}}$/$a_{k}^{\rm{inf}}$ in (22) and (23) is replaced by $\pi^{*}$/$S_{k}$/$a_{k}$ and $\tilde{\pi}^{\rm ip*}$/$\tilde{S}_{k}^{\rm ip}$/$a_{k}^{\rm ip}$ for VoMI and VoII, respectively, while $\pi^{\rm{sup}}$/$S_{k}^{\rm{sup}}$ in (22) and (23) is replaced by $\tilde{\pi}^{*}$/$\tilde{S}_{k}$.
VoMI VoII
EVoI EVoMI EVoII
$\bar{\Xi}=J_{\pi^{*}}-J_{\tilde{\pi}^{*}}$ $\bar{\Xi}=J_{\tilde{\pi}^{\rm ip*}}-J_{\tilde{\pi}^{*}}$
IVoI IVoMI IVoII
$\xi_{k}=A_{\tilde{\pi}^{*}}(\tilde{S}_{k},a_{k})|_{a_{k}\sim\pi^{*}(\cdotp|S_{k})}$ $\xi_{k}=A_{\tilde{\pi}^{*}}(\tilde{S}_{k},a_{k}^{\rm ip})|_{a_{k}^{\rm ip}\sim\tilde{\pi}^{\rm ip*}(\cdotp|\tilde{S}_{k}^{\rm ip})}$
TABLE I: Four categories of utility-based VoI
Case 6.
Consider the vehicle-following control problem in Case 1. The control agent of vehicle $i$ can observe its position error $e_{pi,k}$, velocity error $e_{vi,k}$, and acceleration $acc_{i,k}$ based on local sensory information. Meanwhile, the acceleration $acc_{i-1,k}$ of the predecessor $i-1$ is remote sensory information that has to be transmitted from vehicle $i-1$ to vehicle $i$. In order to quantify the benefit of making decisions when $acc_{i-1,k}$ is available to the control agent of vehicle $i$, the four utility-based VoI can be defined for $acc_{i-1,k}$:
  • EVoMI $\Xi_{i}^{\rm M}=J_{\pi_{i}^{*}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}^{*}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the optimal policies $\pi_{i}^{*}$ and $\tilde{\pi}_{i}^{*}$ for the SSDP in Case 2 and augmented-state SSDP in Case 4, respectively.
  • IVoMI $\xi_{i,k}^{\rm M}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\pi_{i}^{*}(\cdot|S_{i,k})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the augmented-state SSDP in Case 4, and $u_{i,k}\sim\pi_{i}^{*}(\cdot|S_{i,k})$ is the control input of vehicle $i$ that is sampled according to the optimal policy $\pi_{i}^{*}(\cdot|S_{i,k})$ of the SSDP in Case 2.
  • EVoII $\Xi_{i}^{\rm I}=J_{\pi_{i}^{\rm d*}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}^{\rm d*}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the optimal policies $\pi_{i}^{\rm d*}$ and $\tilde{\pi}_{i}^{*}$ for the Random Delay SSDP in Case 5 and augmented-state SSDP in Case 4, respectively.
  • IVoII $\xi_{i,k}^{\rm I}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\tilde{\pi}_{i}^{\rm d*}(\cdot|\tilde{S}_{i,k}^{\rm d})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the augmented-state SSDP in Case 4, and $u_{i,k}\sim\tilde{\pi}_{i}^{\rm d*}(\cdot|\tilde{S}_{i,k}^{\rm d})$ is the control input of vehicle $i$ that is sampled according to the optimal policy $\tilde{\pi}_{i}^{\rm d*}(\cdot|\tilde{S}_{i,k}^{\rm d})$ of the Random Delay SSDP in Case 5.
V-B Information-Theory-based VoI
According to Lemma 1, given the SSDP in Definition 1, a piece of information $I_{k}$ can improve decision making only if it is part of the exogenous information $W_{k}$ or helps in predicting $W_{k}$. Furthermore, as per Definition 1, $W_{k}$ must either affect the transition of state $S_{k}$ and/or the reward $R_{k+1}$. Thus, it follows intuitively that the value of $I_{k}$ is related to its impact on both the state transition and the reward. The VoI can therefore be defined by quantifying, ”How much better could we predict the next state $S_{k+1}$ or the reward $R_{k+1}$ if we included $I_{k}$ in the augmented-state $\tilde{S}_{k}$, compared to if we did not?”
Definition 9.
(Information-Theory-based VoI (ITVoI)) For the SSDP where the state $S_{k}$ does not include information $I_{k}$ and the augmented-state SSDP where the augmented state $\tilde{S}_{k}=\{S_{k},I_{k}\}$ includes $I_{k}$, we convert the state transition functions, the exogenous information function for $I_{k}$, and the reward function, i.e., $f^{S}$, $f^{\tilde{S}}$, $f^{I}$, and $r$ to the corresponding transition probabilities $T^{S}=p\{S_{k+1}|S_{k},a_{k}\}$, $T^{\tilde{S}}=p\{S_{k+1},I_{k+1}|S_{k},a_{k},I_{k}\}$, and $T^{I}=p\{I_{k+1}|I_{k},S_{k},a_{k}\}$, $T^{r(S)}=p\{R_{k+1}|S_{k},a_{k}\}$, and $T^{r(\tilde{S})}=p\{R_{k+1}|S_{k},I_{k},a_{k}\}$. The ITVoI is defined as the sum of the conditional KL divergence of $T^{S}\otimes T^{I}$ from $T^{\tilde{S}}$ and the conditional KL divergence of $T^{r(S)}$ from $T^{r(\tilde{S})}$ as
$ITVoI=D_{KL}(T^{\tilde{S}}||T^{S}\otimes T^{I})+D_{KL}(T^{r(\tilde{S})}||T^{r(S)}),$ (25)
where
$D_{KL}(T^{\tilde{S}}||T^{S}\otimes T^{I})=\int_{\tilde{S}_{k+1},\tilde{S}_{k},a_{k}}p\{\tilde{S}_{k+1},\tilde{S}_{k},a_{k}\}$
$\log\left(\frac{p\{\tilde{S}_{k+1}|\tilde{S}_{k},a_{k}\}}{p\{S_{k+1}|S_{k},a_{k}\}p\{I_{k+1}|S_{k},a_{k},I_{k}\}}\right),$ (26)
and
$D_{KL}(T^{r(\tilde{S})}||T^{r(S)})=\int_{R_{k+1},\tilde{S}_{k},a_{k}}p\{R_{k+1},\tilde{S}_{k},a_{k}\}$
$\log\left(\frac{p\{R_{k+1}|\tilde{S}_{k},a_{k}\}}{p\{R_{k+1}|S_{k},a_{k}\}}\right).$ (27)
Note that the KL divergence in (9) and (9) is a measure of the information lost when $T^{S}\otimes T^{I}$ and $T^{r(S)}$ are used for approximating $T^{\tilde{S}}$ and $T^{r(\tilde{S})}$, respectively. The ITVoI is 00 if the transition of $S_{k}$ and the reward $R_{k+1}$ are independent of $I_{k}$. In this case, we know from Lemma 1 that the performance of the optimal policies of the augmented-state SSDP and the original SSDP are the same, and there is no need to include $I_{k}$ in $\tilde{S}_{k}$. On the other hand, a higher KL divergence value indicates that the transition of $S_{k}$ or the reward $R_{k+1}$ depend on $I_{k}$ to a larger extent, and thus the availability of $I_{k}$ is more important for accurately predicting the future state $S_{k+1}$ or predicting the reward $R_{k+1}$. In this case, including $I_{k}$ in $\tilde{S}_{k}$ will be more likely to improve the performance of the optimal policy. Therefore, the KL divergence is a suitable quantitative measure for the VoI.
Case 7.
Consider the vehicle-following control problem in Case 1. In order to quantify the benefit of making decisions when $acc_{i-1,k}$ is available to the control agent of vehicle $i$, the ITVoI can be calculated for $acc_{i-1,k}$ as
$ITVoI$ $=D_{KL}(T^{\tilde{S}_{i}}||T^{S_{i}}\otimes T^{W_{i}})+D_{KL}(T^{r(\tilde{S}_{i})}||T^{r(S_{i})})$
$=\int_{S_{i,k},acc_{i-1,k},u_{i,k}}p\{S_{i,k},acc_{i-1,k},u_{i,k}\}$
$\log\left(\frac{1}{p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}}\right).$ (28)
From (7), we can see that the ITVoI of $acc_{i-1,k}$ depends on the conditional probability $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}$, i.e., how well we can predict the acceleration $acc_{i-1,k}$ of predecessor $i-1$ given the driving status $S_{i,k}$ and control input $u_{i,k}$ of the following vehicle $i$? Consider the extreme case when $acc_{i-1,k}=g(S_{i,k},u_{i,k})$ is a deterministic function of $S_{i,k}$ and/or $u_{i,k}$, we thus have $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}=\mathbf{1}_{acc_{i-1,k}=g(S_{i,k},u_{i,k})}$. Here, the ITVoI of $acc_{i-1,k}$ is 00, meaning there is no value in knowing $acc_{i-1,k}$. Note that this scenario is not possible in practice, since $acc_{i-1,k}$ is a function of $acc_{i-1,k-1}$ and $u_{i-1,k-1}$ according to (2). The predecessor $i-1$ cannot determine its control input $u_{i-1,k-1}$ at control interval $k-1$ based on $S_{i,k}$ and/or $u_{i,k}$, whose values are not available until control interval $k$. In the other extreme case, if the predecessor $i-1$ makes driving decisions completely independently of the following vehicle $i$, we have $p\{acc_{i-1,k}|S_{i,k},u_{i,k}\}=p\{acc_{i-1,k}\}$. This means the ITVoI depends on the probability distribution of $acc_{i-1,k}$. For example, if $acc_{i-1,k}$ is a constant value, i.e., $p\{acc_{i-1,k}=c\}=1$, the ITVoI is 00.
Compared to the utility-based VoI, the ITVoI is generally easier to calculate when the model of the environment, i.e., the state transition probabilities and the reward function, is known. This advantage arises because ITVoI does not require deriving the optimal policies of the SSDPs with state $S_{k}$ and augmented state $\tilde{S}_{k}$. However, there are several significant limitations to ITVoI. Firstly, it requires the knowledge of the model of the environment, which might not be accurately available in many practical tasks. Secondly, among the four types of utility-based VoI, the ITVoI is essentially equivalent to the EVoMI, as both aim to predict the benefit of having additional information $I_{k}$ before making decisions in terms of long-term average performance. However, ITVoI does not cover the other three types of utility-based VoI, i.e., IVoMI, EVoII, and IVoII, thereby limiting its applicability. Finally, the utility-based VoI captures the ultimate objective of information usage, which is to improve the control performance. Meanwhile, the ITVoI measures the merit of information for improving some intermediate outcomes that could contribute to the enhancement of the ultimate objective. Therefore, utility-based VoI is more intuitive compared to ITVoI. In the next section, we will focus on the estimation of utility-based VoI.
VI VoI Estimation using DRL
To obtain either EVoI or IVoI, we have to derive both policies $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}_{k}$ in Definition 7 and Definition 8. This can be achieved by using the DRL algorithms to train the control agent. The inferior policy $\pi^{\rm{inf}}$ is learned in the environment with partial/imperfect observations $S^{\rm{inf}}_{k}$, while the superior policy $\pi^{\rm{sup}}$ is learned in the environment with full/perfect observations $S^{\rm{sup}}_{k}$.
VI-A EVoI Estimation
After training the two policies, EVoI can be derived by testing the two policies for multiple episodes, deriving the performance of each policy in terms of the average cumulative rewards, and finally calculating the performance difference between the two policies.
VI-B IVoI Estimation
The IVoI corresponds to the advantage function of the superior policy $\pi^{\rm{sup}}$ with full/perfect observations $S^{\rm{sup}}_{k}$, where various methods can be used to estimate the advantage function of a policy in RL.
VI-B1 Method A
By the definition of advantage function in (14), we can train two neural networks (NNs) to represent the value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$ and Q function $q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},a_{k})$, respectively. The IVoI can be derived simply as their difference
$\xi_{k}$ $=A_{\pi^{\rm{sup}}}(S_{k}^{\rm{sup}},a_{k}^{\rm{inf}})|_{a_{k}^{\rm{inf}}\sim\pi^{\rm{inf}}(\cdotp|S_{k}^{\rm{inf}})}$
$=q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},a_{k}^{\rm{inf}})|_{a_{k}^{\rm{inf}}\sim\pi^{\rm{inf}}(\cdotp|S_{k}^{\rm{inf}})}-v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k}).$ (29)
However, it is a little cumbersome to train two NNs. Moreover, if the NN that represents the Q function $q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},a_{k})$ is trained during learning of the superior policy $\pi^{\rm{sup}}$, it might provide a less accurate estimation of the Q values for those actions that are favored by the inferior policy $\pi^{\rm{inf}}$ but seldomly chosen by $\pi^{\rm{sup}}$.
VI-B2 Method B
To solve the above issues, another method is to only train one NN to represent the value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$, and estimate the advantage function using the Temporal Difference (TD) error
$\delta_{\pi^{\rm{inf}},k}=R_{k+1}+\gamma v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k+1})-v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k}).$ (30)
In this case, the IVoI can be estimated for each control interval based on the reward $R_{k+1}$ as a result of implementing action $a_{k}^{\rm{inf}}\sim(a_{k}|S^{\rm{inf}}_{k})$, as well as the value function $v_{\pi^{\rm{sup}}}$ of the optimal control policy for the current state $S^{\rm{sup}}_{k}$ and the next state $S^{\rm{sup}}_{k+1}$. Since we have
$\xi_{k}$ $=A_{\pi^{\rm{sup}}}(S_{k}^{\rm{sup}},a_{k}^{\rm{inf}})|_{a_{k}^{\rm{inf}}\sim\pi^{\rm{inf}}(\cdotp|S_{k}^{\rm{inf}})}$
$=\mathrm{E}_{S^{\rm{sup}}_{k+1}}[\delta_{\pi^{\rm{inf}},k}],$ (31)
$\delta_{\pi^{\rm{inf}},k}$ is an unbiased estimate of the IVoI if the NN can accurately represent the true value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$. This method of estimation is subject to noise. However, the variance remains relatively low, as the expectation in (VI-B2) is taken only with respect to the next state $S^{\rm{sup}}_{k+1}$.
VII VoI-Oriented Optimization in Communication Systems
When full and perfect information corresponding to the environment state is available, the control agent can typically make decisions that yield better performance than those based on partial and/or imperfect observations. However, radio resources in wireless networks are limited and valuable. Therefore, the communication agents have to make decisions on information transmission and RRM in order to strike a good tradeoff between minimizing control performance degradation due to non-ideal communication and maximizing communication performance for traditional services.
VII-A Three Types of Communication Decisions
The communication decisions can be divided into three broad categories as explained below.
VII-A1 What to Communicate
The remote sensory information normally consists of multiple elements. The communication agents can make decisions on
  • the subset of information that should be transmitted;
  • the processed information for transmission.
VII-A2 When to Communicate
It might not be necessary to transmit a piece of (processed) information at every control interval, depending on the specific instance of information. Therefore, the communication agent can make decisions on whether the element of information should be transmitted or not at each control interval based on its content and relevance with the previous information.
VII-A3 How to Communicate
For a piece of (processed) information to be transmitted at a control interval, the radio resources need to be scheduled to support its transmission. Considering the stringent resource limitations, the communication agents can make decisions to schedule available resources for transmission of the most meritorious information.
VII-B Three Time Scales for Communication Decision-Making
VII-B1 Every Communication Interval - How to Communicate
Since RRM decisions are normally made at a faster time scale (e.g., every $1$ ms in fourth-generation (4G) cellular system) than the control decisions, one control interval can be further divided into $T$ communication intervals. Let $\Delta t$ represent the duration of a communication interval. We denote the $t$-th communication interval in control interval $k$ as communication interval $(k,t)$, where $t\in\mathcal{T}=\{0,1,2,\cdots,T-1\}$. The “How to Communicate” decisions are made at every communication interval.
VII-B2 Every Control Interval - When and What to Communicate
Since sensory information is sampled at the beginning of each control interval, the “When to Communicate” and “What to Communicate” decisions can be made dynamically per control interval. The communication agent determines whether to communicate with its peers and, if so, selects the subset of information or processed information to share during this control interval.
VII-B3 Across Multiple Control Intervals - When and What to Communicate
The “When to Communicate” and “What to Communicate” decisions can also be made statically only once at control interval 00 or semi-statically for every $C>1$ control intervals. In this case, the subset of information or processed information will always be transmitted at every control interval within the effective time range of the decision. An example is choosing the information topology for a platoon [28].
Case 8.
We consider a typical urban C-V2X network, where V2V links coexist with V2I links. A V2I link connects a vehicle to the BS and can be used for high-throughput services. A V2V link connects a pair of preceding vehicle $i-1$ and following vehicle $i$ for periodic transmission of cooperative awareness messages (CAM) to support advanced driving services. Specifically, at the beginning of each control interval $k$, the acceleration $acc_{i-1,k}$ of the predecessor $i-1$ is sampled to form the CAM and buffered in a queue before transmission to the corresponding follower $i$.
At each communication interval $(k,t)$, the predecessor $i-1$ transmits the CAM generated at the beginning of each control interval $k$. Any CAM that is not fully transmitted by the end of the last communication interval $(k,T-1)$ of control interval $k$ will be discarded. In other words, any CAM not fully delivered during the previous control interval $k-1$ is replaced by newly generated CAM at the beginning of control interval $k$.
As introduced in Case 1, the control agent in follower $i$ determines its control input based on the local sensory information $ep_{i,k}$, $ev_{i,k}$, $acc_{i,k}$, as well as the remote sensory information received from its predecessor $i-1$.
There is a communication agent in each predecessor $i-1$ to make communication decisions, which fall into one or more of the following categories:
  • Every Communication Interval - How to Communicate: The communication agent makes RRM decisions at each communication interval $(k,t)$. In order to enhance spectrum utilization, we consider that $L$ V2V links share the uplink resources orthogonally allocated to $M$ V2I links. Without loss of generality, we assume that every V2I link $m\in\mathcal{M}=\{0,\cdots,M-1\}$ is pre-assigned sub-channel $m$ with constant transmit power $P^{\rm I}_{m}$. Moreover, one or more V2V links $i-1\in\mathcal{L}=\{1,3,5\cdots,2L-1\}$ can reuse the sub-channels of the V2I links for CAM transmission. We use the binary allocation indicator $\theta_{i-1,m,(k,t)}\in\{0,1\}$ to indicate whether V2V link $i-1$ occupies sub-channel $m$ at communication interval $(k,t)$ or not. Moreover, we consider that each V2V link $i-1$ occupies at most one sub-channel, i.e., $\sum_{m=0}^{M-1}\theta_{i-1,m,(k,t)}\leq 1$. Let $P^{\rm V}_{i-1,m,(k,t)}$ denote the transmit power of V2V link $i-1$ over the sub-channel $m$ at communication interval $(k,t)$. Therefore, each communication agent $i-1$ has to make the subchannel allocation $\{\theta_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ and power allocation $\{P^{\rm V}_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ decisions at each communication interval $(k,t)$.
  • Every Control Interval - When and What to Communicate: Let $\phi_{i-1,k}\in\{0,1\}$ be the binary transmission indicator to indicate whether the sampled CAM $acc_{i-1,k}$ will be transmitted to the follower $i$ or not. At each control interval $k$, each communication agent $i-1$ can make decisions on $\phi_{i-1,k}$.
  • Across Multiple Control Intervals - When and What to Communicate: Each communication agent $i-1$ makes a one-time decision on whether it will transmit its acceleration or not to its follower $i$. Once the decision is made, the predecessor will always transmit $acc_{i-1,k}$ or not at each control interval $k$.
VII-C Optimization Objective
In general, the optimization objective is to maximize communication performance for traditional services, while minimizing the control performance degradation in vehicle control due to non-ideal communications, i.e.,
$\max_{\pi_{\rm CM}}J^{\mathrm{CM}}=\max_{\pi_{\rm CM}}(\kappa_{1}J^{\mathrm{CM\_trad}}+\kappa_{2}\bar{\Xi}),$ (32)
where $J^{\mathrm{CM\_trad}}$ represents the long-term average communication performance for traditional services, such as the average throughput, and $\bar{\Xi}$ is the EVoI. The weighting coefficients $\kappa_{1}$ and $\kappa_{2}$ reflect the relative importance of traditional network services and control services in the optimization objective.
According to Lemma 2, the optimization objective in (32) can be expressed in the following form of expectation.
$J^{\mathrm{CM}}$
$=\mathrm{E}_{\pi^{\rm CM}}\sum_{k=0}^{\infty}\left[\kappa_{1}\left(\sum_{t=0}^{T-1}\gamma^{kT+t}r_{(k,t)}^{\mathrm{CM\_trad}}\right)+\kappa_{2}\gamma^{k}\mathrm{E}_{\pi^{\rm inf}}[\xi_{k+1}]\right],$ (33)
where $r_{(k,t)}^{\mathrm{CM\_trad}}$ represents the instantaneous communication performance at communication interval $(k,t)$, e.g., instantaneous data rate, where
$J^{\mathrm{CM\_trad}}=\mathrm{E}_{\pi^{\rm CM}}\left[\sum_{k=0}^{\infty}\sum_{t=0}^{T-1}\gamma^{kT+t}r_{(k,t)}^{\mathrm{CM\_trad}}\right],$ (34)
and $\xi_{k+1}$ is the IVoI.
Case 9.
In Case 8, the optimization objective of the communication agents $i-1\in\mathcal{L}$ is to maximize the sum V2I throughtput while minimizing the performance degradation of the followers $i$ due to non-ideal communications.
The instantaneous channel gain of V2V link $i-1$ over sub-channel $m$ (occupied by V2I link $m$) at communication interval $(k,t)$ is denoted by $G_{i-1,m,(k,t)}$. $G_{i-1,m,(k,t)}$ remains constant within a communication interval, and can be written as
$G_{i-1,m,(k,t)}=\alpha_{i-1,(k,t)}h_{i-1,m,(k,t)},$ (35)
where $\alpha_{i-1,(k,t)}$ captures the large-scale fading effects including path loss and shadowing, which are assumed to be frequency independent. $h_{i-1,m,(k,t)}$ corresponds to frequency-dependent small-scale fading.
Similarly, let $G_{m,(k,t)}$ denote the instantaneous channel gain of the V2I link $m$; $G_{i-1,B,m,(k,t)}$ the interference channel gain from V2V link $i-1$ transmitter at predecessor $i-1$ to V2I link $m$ receiver at the BS; $G_{B,i-1,m,(k,t)}$ the interference channel gain from V2I link $m$ transmitter at the BS to V2V link $i-1$ receiver at follower $i$; and $G_{j-1,i-1,m,(k,t)}$ the interference channel gain from the V2V link $j-1$ transmitter at predecessor $j-1$ to the V2V link $i-1$ receiver at follower $i$ over the sub-channel $m$.
The SINR $\gamma_{m,(k,t)}$ of V2I link $m$ and the SINR $\gamma_{i-1,m,(k,t)}$ of V2V link $i-1$ on sub-channel $m$ at communication interval $(k,t)$ are derived by
$\gamma_{m,(k,t)}=\frac{P^{\rm I}_{m}G_{m,(k,t)}}{\sigma^{2}+\sum\limits_{i-1\in\mathcal{L}}\phi_{i-1,k}\theta_{i-1,m,(k,t)}P^{\rm V}_{i-1,m,(k,t)}G_{i-1,B,m,(k,t)}},$ (36)
and
$\gamma_{i-1,m,(k,t)}=\frac{P^{\rm V}_{i-1,m,(k,t)}G_{i-1,m,(k,t)}}{\sigma^{2}+I_{i-1,m,(k,t)}},$ (37)
respectively. $\sigma^{2}$ is the power of channel noise which satisfies the independent Gaussian distribution with a zero mean value. $I_{i-1,m,(k,t)}$ is the total interference power received by V2V link $i-1$ over sub-channel $m$, where
$I_{i-1,m,(k,t)}=$
$P^{\rm I}_{m}G_{B,i-1,m,(k,t)}+$
$\sum\limits_{j-1\in\mathcal{L}\backslash\{i-1\}}\phi_{i-1,k}\theta_{j-1,m,(k,t)}P^{\rm V}_{j-1,m,(k,t)}G_{j-1,i-1,m,(k,t)}.$
The instantaneous data rates $r_{m,(k,t)}$ and $r_{i-1,(k,t)}$ of V2I link $m$ and V2V link $i-1$ at communication interval $(k,t)$ are respectively derived as
$r_{m,(k,t)}=B\log_{2}(1+\gamma_{m,(k,t)}),$ (38)
and
$r_{i-1,(k,t)}=\sum_{m=0}^{M-1}{\theta_{i-1,m,(k,t)}}B\log_{2}(1+\gamma_{i-1,m,(k,t)}),$ (39)
where $B$ is the bandwidth of a sub-channel.
Let $r^{\rm CAM}_{i-1,(k,t)}$ denote the transmission rate of V2V link $i-1$ in terms of CAM at communication interval $(k,t)$, which is given by
$r^{\rm CAM}_{i-1,(k,t)}=\frac{r_{i-1,(k,t)}}{N_{c}},$ (40)
where $N_{c}$ is the constant CAM size.
Each predecessor $i-1\in\mathcal{L}$ has a buffer to store its CAM. Let $q^{\rm CAM}_{i-1,(k,t)}$ denote the queue length of predecessor $i-1$, in terms of the number of CAMs, at communication interval $(k,t)$. At the beginning of the first communication interval $(k,0)$ of each control interval $k\in\mathcal{K}$, predecessor $i-1$ samples the driving status for control interval $k$ and buffers the generated CAM. At each communication interval $(k,t),t\in\mathcal{T}$, within control interval $k$, the queue length $q^{\rm CAM}_{i-1,(k,t)}$ is decreased by $\Delta t\times r^{\rm CAM}_{i-1,(k,t)}$, which is the number of CAM transmitted during the communication interval. Since any CAM that is not fully transmitted by the end of the last communication interval $(k,T-1)$ of control interval $k$ will be discarded, at the start of each new control interval $k\in\mathcal{K}$, specifically at communication interval $(k,0)$, $q^{\rm CAM}_{i-1,(k,0)}$ is set to $1$. Since the value of $q^{\rm CAM}_{i-1,(k,0)}$ changes abruptly at the beginning of each control interval $k$, we use $q^{\rm CAM}_{i-1,(k,T)}$ to denote the queue length by the end of the last communication interval $(k,T-1)$ of control interval $k$. Note that $q^{\rm CAM}_{i-1,(k,T)}\neq q^{\rm CAM}_{i,(k+1,0)}$ since the queue length is discontinuous at the control interval boundaries.
The queue process evolves as
$q^{\rm CAM}_{i-1,(k,t)}=$
$\left\{\begin{array}[]{ll}1,&\mathrm{if}\quad t=0\\ \max\left[0,q^{\rm CAM}_{i-1,(k,t-1)}-\Delta t\times r^{\rm CAM}_{i-1,(k,t-1)}\right],&\mathrm{if}\quad 0<t\leq T\\ \end{array}\right..$ (43)
For each follower $i$, if a CAM is fully delivered to it by the end of control interval $k$, the receiver will replace its old CAM with the newly received CAM at the beginning of control interval $k+1$. If a CAM is not fully delivered by the end of the control interval, the receiver will continue to use the last fully received CAM, as it cannot interpret an incomplete CAM. When the most recent available CAM of follower $i$ at the beginning of control interval $k$ corresponds to the CAM sampled at the predecessor $i-1$ in control interval $k-\tau_{i,k}$, we refer to $\tau_{i,k}$ as the observation delay at V2V receiver $i$, which is derived from the queue length as
$\tau_{i,k}=\left\{\begin{array}[]{ll}1,&\mathrm{if}\quad q^{\rm CAM}_{i-1,(k,T)}=0\\ \tau_{i,k}+1,&\mathrm{otherwise}\\ \end{array}\right..$ (46)
Notably, each follower $i$ makes decisions based on the augmented state of a Random Delay SSDP as given in Case 5.
The optimization objective is formulated as
$J^{\mathrm{CM}}=\kappa_{1}\sum_{m=0}^{M-1}\mathrm{E}_{\pi^{\rm CM}}\left[\sum_{k=0}^{K-1}\sum_{t=0}^{T-1}\gamma^{kT+t}r_{m,(k,t)}\right]+\kappa_{2}\sum_{i-1\in\mathcal{L}}\Xi_{i},$ (47)
where $\gamma$ is the RRM reward discount factor. The first term corresponds to the discounted sum throughput of all V2I links $m\in\mathcal{M}$ over all the communication intervals of a control episode, and the second term corresponds to the EVoMI $\Xi_{i}^{\rm M}$ or EVoII $\Xi_{i}^{\rm I}$ of follower $i$ defined in Case 6. The weight factors $\kappa_{1}$ and $\kappa_{2}$ indicate the relative importance of minimizing control performance degradation versus maximizing the V2I throughput.
The optimization objective $J^{\mathrm{CM}}$ should correspond to the expected return of the DRL model, i.e.,
$J^{\mathrm{CM}}=\mathrm{E}_{\pi^{\rm CM}}\mathrm{E}_{\pi^{\rm inf}}\left[\sum_{k=0}^{K-1}\sum_{t=0}^{T-1}\gamma^{kT+t}r^{\rm CM}_{(k,t)}\right],0\leq\eta\leq 1,$ (48)
where the expectation is taken with respect to the probability distribution of the state-action trajectories when the RRM policy is $\pi^{\rm CM}$ and vehicle control policy is $\pi^{\mathrm{inf}}$.
VII-D VoI-oriented Optimization for Vehicular Communication
When addressing the problem of ”When and What to Communicate”, different approaches might be adopted depending on the decision-making time scale, i.e.,
  • Across Multiple Control Intervals: When decisions are made statically or semi-statically across multiple control intervals, we can evaluate the $J^{\mathrm{CM}}$ for different decisions and choose the decision that maximizes $J^{\mathrm{CM}}$. To evaluate $J^{\mathrm{CM}}$, the EVoI can be estimated as outlined in Section V, and $J^{\mathrm{CM\_trad}}$ can be evaluated using theoretical tools such as queuing theory or through simulation.
  • Every Control Interval: When decisions are made dynamically per control interval, we can formulate an SSDP to optimize the communication decisions made at each control interval. Specifically, the reward function of the SSDP can be written as
    $r^{\rm CM}_{(k,t)}=\kappa_{1}\left(\sum_{t=0}^{T-1}\gamma^{kT+t}r_{(k,t)}^{\mathrm{CM\_trad}}\right)+\kappa_{2}\gamma^{k}\xi_{k+1}.$ (49)
    Note that $\xi_{k+1}=A_{\pi^{\text{sup}}}(S_{k+1}^{\text{sup}},a_{k+1})|_{a_{k+1}^{\text{inf}}\sim\pi^{\text{inf}}(\cdot|S_{k+1}^{\text{inf}})}$ represents the amount of control performance loss that the information recipient will suffer from, as a consequence of its control agent having to make a decision at the next control interval $k+1$ based on the partial/imperfect observation $S_{k+1}^{\text{inf}}$ according to the inferior policy $\pi^{\text{inf}}(\cdot|S_{k+1}^{\text{inf}})$ instead of based on the full/perfect observation $S_{k+1}^{\text{sup}}$ according to the superior policy $\pi^{\text{sup}}(\cdot|S_{k+1}^{\text{sup}})$. The partial/imperfect observation received at control interval $k+1$ is affected by the “When and What to Communicate” decisions made at the beginning of control interval $k$ by the communication agent for the information source. When $\xi_{k+1}$ in (49) corresponds to the IVoMI of a piece of information, we can determine whether the instance of information generated at control interval $k$ will be transmitted or not by the information source. When $\xi_{k+1}$ corresponds to the IVoII of a piece of information, we can determine what processed information should be shared by the information source.
For the “How to Communicate” Problem, we can formulate an SSDP to optimize the communication decisions made at each communication interval. Specifically, the reward function of the SSDP can be written as
$r^{\rm CM}_{(k,t)}=\left\{\begin{array}[]{ll}\kappa_{1}r_{(k,t)}^{\mathrm{CM\_trad}},&\ 0\leq t<T-1\\ \kappa_{1}r_{(k,t)}^{\mathrm{CM\_trad}}+\kappa_{2}\xi_{k+1},&\ t=T-1\\ \end{array}\right..$ (52)
Note that the $r_{(k,t)}^{\mathrm{CM\_trad}}$ component is present in the reward of every communication interval, while the IVoII component, i.e., $\xi_{k+1}$, is only present in the last communication interval of each control interval when $t=T-1$. The imperfect observation received by the control agent at control interval $k+1$ is affected by the degree of communication nonideality the transmitted data has experienced, which is in part a result of the sequence of decisions on communication actions $a_{(k,0)}^{\text{CM}},a_{(k,1)}^{\text{CM}},\ldots,a_{(k,T-1)}^{\text{CM}}$ that are made during control interval $k$ by the communication agent.
Since EVoI equals the expected cumulative IVoI according to Lemma 2, the expected discounted sum of the immediate rewards in (49) and (52) equal the optimization objective in (32).
Case 10.
The state space, action space, and reward function of the SSDPs for making communication decisions in Case 9 are given below. The SSDPs can be solved by a centralized communication agent at the BS; or by the decentralized communication agents at each predecessor $i-1\in\mathcal{L}$. The algorithms to solve the SSDP models are out of the scope of this paper. We refer the readers to [27] for more details on a similar example.
  • State Space
    • When and What to Communicate:
      $S^{\rm CM}_{(k)}=\{\bm{G}_{(k,0)},\{\tau_{i,k}\}_{i-1\in\mathcal{L}},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}+1}^{k}\}$
    • How to Communicate:
      $S^{\rm CM}_{(k,t)}=\{\bm{G}_{(k,t)},q^{\rm CAM}_{(k,t)},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}+1}^{k},t\}$,
  • Action Space
    • When and What to Communicate:
      $a^{\rm CM}=\{\phi_{i-1,k}\}_{m\in\mathcal{M},i-1\in\mathcal{L}}$
    • How to Communicate:
      $a^{\rm CM}=\{\theta_{i-1,m,(k,t)},P^{\rm V}_{i-1,m,(k,t)}\}_{m\in\mathcal{M},i-1\in\mathcal{L}}$.
  • Reward Function
    • When and What to Communicate:
      $r^{\rm CM}_{(k,t)}$ $=\kappa_{1}\left(\sum_{m=0}^{M-1}\sum_{t=0}^{T-1}\gamma^{kT+t}r_{m,(k,t)}\right)$
      $+\kappa_{2}\sum_{i-1\in\mathcal{L}}\gamma^{k}\xi_{i,k+1}.$ (53)
    • How to Communicate:
      $r^{\rm CM}_{(k,t)}=$
      $\left\{\begin{array}[]{ll}\kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)},&\ 0\leq t<T-1\\ \kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)}+&\\ \kappa_{2}\sum_{i-1\in\mathcal{L}}\xi_{i,k+1}^{\rm I},&\ t=T-1\\ \end{array}\right..$ (57)
VIII Conclusion and Future Work
The significance of this paper is to establish a canonical framework for modeling the value of the information shared over 6G V2X networks, specifically for vehicle control decisions under various forms of uncertainty. To lay the theoretical groundwork, we have presented the SSDP model as an extension of the MDP model, explicitly accounting for useful information that may be missing from the state. This extension accommodates both fully observable and partially observable scenarios. Based on the SSDP model, we have proposed a mathematically rigorous and practical method to identify the set of information that can enhance decision-making if made available.
At the heart of this framework are two interrelated definitions of VoI — EVoI and IVoI — derived from the SSDP model. These definitions correspond to the regret of a policy with partial or imperfect observations and the advantage function of a policy with full or perfect observations, respectively, in the context of RL terminology.
The primary purpose of VoI modeling is to enable the integrated design and joint optimization of control and communication. Since VoI should be estimated by the control agent of the information recipient and shared with the communication agent of the information source to guide optimization of communication decisions, we have provided methods for estimating VoI using DRL, and for formulating the SSDP model to optimize communication decisions based on the estimated VoI. We have used a simple vehicle-following control problem as an illustrative example throughout the paper.
While the proposed framework holds great promise for enabling the joint optimization of stochastic, sequential control and communication decisions across diverse networked control systems, several key challenges present compelling directions for future research.
Firstly, accurate VoI estimation by control agents is crucial for providing precise reward signals that guide the communication agents’ decision-making optimization. The estimation methods discussed in this paper, particularly those for IVoI, still have room for improvement. This is because IVoI is essentially an advantage function, and existing estimation methods for advantage functions are primarily focus on learning optimal policies rather than accurately estimating the advantage functions themselves. As a result, the precision of the learned values - especially for those states and actions that are infrequently visited under policies with full or perfect observations - may not be sufficient for accurate VoI estimation.
Secondly, the construction of communication state variables in SSDP models plays a critical role in the performance of the learned communication policies. Generally speaking, the communication state should comprise two components. The first component includes Channel State Information (CSI), Queue State Information (QSI), and other elements typically defined in MDP models for communication problems. Meanwhile, the second component pertains to IVoI, which theoretically should include any information that is useful in predicting the IVoI at the recipient. The IVoI, as an advantage function, depends on the recipient’s full or perfect observations, the action determined by the policy under partial or imperfect observations, and the neural network parameters for the critic of the policy with full or perfect observations. However, since the communication agent at the information source cannot obtain or construct the state using all this information, developing a systematic method for communication state construction that is independent of specific control tasks remains an open issue.
Lastly, since the VoI only appears in the reward of the last communication interval within each control interval in (10), the challenge of learning efficient communication policies under a sparse reward setting must be addressed.
-A Proof of Theorem 1
When the SSDP in Definition 1 is an MDP, we have
$\mathrm{Pr}(S_{k+1}|S_{k},a_{k})=\mathrm{Pr}(S_{k+1}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1}).$ (58)
Therefore, from (58), we have
$\mathrm{Pr}(f^{S}(S_{k},a_{k},f^{W}(S_{k},a_{k},\xi_{k}))|S_{k},a_{k})$
$=$ $\mathrm{Pr}(f^{S}(S_{k},a_{k},f^{W}(S_{k},a_{k},\xi_{k}))|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$ (59)
which follows from the system’s state transition function $S_{k+1}=f^{S}(S_{k},a_{k},W_{k})$ given in Definition 1. In addition, $W_{k}=f^{W}(\tilde{W}_{k})$ according to the exogenous information function in Definition 1, while $\tilde{W}_{k}\subseteq\{S_{k},a_{k},\xi_{k}\}$ as stated in Theorem 1.
From (-A), we have
$\mathrm{Pr}(\xi_{k}|S_{k},a_{k})$
$=$ $\mathrm{Pr}(\xi_{k}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1})$
$=$ $\mathrm{Pr}(\xi_{k}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$
where the last equation follows from $W_{k^{\prime}}=f^{\rm WS}(S_{k^{\prime}+1},S_{k^{\prime}},a_{k^{\prime}})$ since $S_{k^{\prime}+1}=f^{\rm S}(S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}})$, and the property of conditional probability $\mathrm{Pr}(X|g(Y),Y)=\mathrm{Pr}(X|Y)$.
Since $W_{k}=f^{W}(S_{k},a_{k},\xi_{k})$, $\xi_{k}$ is a set of random variables whose distribution is independent of $S_{k}$ and $a_{k}$ by the expression of the formula. Thus, from (-A), we have
$\mathrm{Pr}(\xi_{k})=\mathrm{Pr}(\xi_{k}|\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$ (61)
which shows that the distribution of $\xi_{k}$ does not depend on the past states, actions and exogenous information, i.e., $\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}$.
Meanwhile, the reverse process of the proof is also true. This completes the proof of Theorem 1.
-B Proof of Theorem 2
We first prove the following Propostion 1.
Proposition 1.
Given the SSDP in Definition 1 is also an MDP, we have
$\mathrm{Pr}(W_{k}|S_{k},\tilde{S}_{k-1},a_{k-1})=\mathrm{Pr}(W_{k}|S_{k})$ (62)
Given the SSDP in Definition 1 is also an MDP, we have $\mathrm{Pr}(W_{k}|S_{k},a_{k})=\mathrm{Pr}(W_{k}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1})$. Taking the expectation with regard to $a_{k}$ on both sides of the equation, we have
$\mathrm{Pr}(W_{k}|S_{k})$ $=\mathrm{Pr}(W_{k}|S_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1})$
$=\mathrm{Pr}(W_{k}|S_{k},S_{k-1},a_{k-1})$
$=\mathrm{Pr}(W_{k}|S_{k},S_{k-1},a_{k-1},W_{k-1})$
where the last equation follows from $W_{k-1}=f^{\rm WS}(S_{k},S_{k-1},a_{k-1})$ since $S_{k}=f^{\rm S}(S_{k-1},a_{k-1},W_{k-1})$, and the property of conditional probability $\mathrm{Pr}(X|g(Y),Y)=\mathrm{Pr}(X|Y)$. Thus, Proposition 1 is proved.
When the SSDP in Definition 1 is not an MDP, we can first convert it to an MDP by including the history in the state. In the following, we will only consider the case when the SSDP in Definition 1 is an MDP. In this case, the augmented-state SSDP is also an MDP having a state transition function
$\tilde{S}_{k+1}$ $=\begin{pmatrix}S_{k+1}\\ W_{k+1}\end{pmatrix}$
$=\begin{pmatrix}f^{S}(S_{k},a_{k},W_{k})\\ f^{W}(f^{S}(S_{k},a_{k},W_{k}),\tilde{\pi}({f^{S}(S_{k},a_{k},W_{k})}),\xi_{k})\end{pmatrix}$
$=f^{\tilde{S}}(\tilde{S}_{k},a_{k},\xi_{k}),$ (64)
where the exogenous information $\xi_{k}$ is an independent random variable with given distribution.
We first consider the finite horizon problem, where there are $K$ time steps indexed by $k\in\{0,1,\cdots,K-1\}$.
Let $v_{k}^{*}(S_{k})$ and $\tilde{v}_{k}^{*}(\tilde{S}_{k})$ denote the value functions under the optimal policies $\pi^{*}$ and $\tilde{\pi}^{*}$ for the original SSDP and augmented-state SSDP, respectively. Define $\tilde{v}_{k}^{*}(S_{k})=\mathrm{E}_{W_{k}}[\tilde{v}_{k}^{*}(\tilde{S}_{k})|S_{k}]$.
Note that $\tilde{J}^{*}=\mathrm{E}_{\tilde{S}_{0}}[\tilde{v}_{0}^{*}(\tilde{S}_{0})]=\mathrm{E}_{S_{0}}[\tilde{v}_{0}^{*}(S_{0})]$ and $J^{*}=\mathrm{E}_{S_{0}}[v_{0}^{*}(S_{0})]$. Therefore, in order to prove that $\tilde{J}^{*}\geq J^{*}$, it is sufficient to prove
$\tilde{v}_{k}^{*}(S_{k})\geq v_{k}^{*}(S_{k}),\forall\ S_{k}\ \mathrm{and}\ k.$ (65)
We will show (65) by induction. For the last time step $K$, we have
$\tilde{v}_{K}^{*}(S_{K})=\mathrm{E}_{W_{K}}\Big{[}\max_{\tilde{\mu}_{K}(\tilde{S}_{K})}r\big{(}\tilde{S}_{K},\tilde{\mu}_{K}(\tilde{S}_{K})\big{)}|S_{K}\Big{]},\forall S_{K},$ (66)
$v_{K}^{*}(S_{K})=\max_{\mu_{K}(S_{K})}\mathrm{E}_{W_{K}}\Big{[}r\big{(}\tilde{S}_{K},\mu_{K}(S_{K})\big{)}|S_{K}\Big{]},\forall S_{K}.$ (67)
According to (66) and (67), we have
$\tilde{v}_{K}^{*}(S_{K})\geq v_{K}^{*}(S_{K}),\forall S_{K},$ (68)
(since we generally have $\mathrm{E}[\max\{\cdot\}]\geq\max\{E[\cdot]\}$ according to Jensen’s inequality). Therefore, the optimal action $\tilde{\pi}_{K}(\tilde{S}_{K})$ for the augmented-state SSDP problem is at least as good as that for the original SSDP $\pi_{K}(S_{K})$ at time step $K$.
Assume that
$\tilde{v}_{k+1}^{*}(S_{k+1})\geq v_{k+1}^{*}(S_{k+1}),\forall S_{k+1}.$ (69)
Consider the Bellman Equation for the original SSDP as
$v_{k}^{*}(S_{k})=\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\biggl{[}r\bigl{(}S_{k},\pi_{k}(S_{k}),W_{k}\bigr{)}+v_{k+1}^{*}\bigl{(}f^{S}(S_{k},$
$\pi_{k}(S_{k}),W_{k})\bigr{)}\Big{|}S_{k},\pi_{k}(S_{k})\biggr{]}\Biggr{\}},$ (70)
and consider the Bellman Equation for the augmented-state problem as
$\tilde{v}_{k}^{*}(\tilde{S}_{k})=\max_{\tilde{\pi}_{k}(\tilde{S}_{k})}\Biggl{\{}r\bigl{(}\tilde{S}_{k},\tilde{\pi}_{k}(\tilde{S}_{k})\bigr{)}+\mathrm{E}_{\tilde{S}_{k+1}}\biggl{[}\tilde{v}_{k+1}^{*}(S_{k+1},$
$W_{k+1})\Big{|}\tilde{S}_{k},\tilde{\pi}_{k}(\tilde{S}_{k})\biggr{]}\Biggr{\}}.$ (71)
Taking the expectation over $W_{k}$ conditioned on $S_{k}$ on both sides of (-B), we have the following Bellman equation
$\tilde{v}_{k}^{*}(S_{k})=\mathrm{E}_{W_{k}}\bigl{[}\tilde{v}_{k}^{*}(\tilde{S}_{k})|S_{k}\bigr{]}$
$=\mathrm{E}_{W_{k}}\Biggl{[}\max_{\tilde{\pi}_{k}(\tilde{S}_{k})}\Biggl{\{}r\bigl{(}\tilde{S}_{k},\tilde{\pi}_{k}(\tilde{S}_{k})\bigr{)}+\mathrm{E}_{\tilde{S}_{k+1}}\biggl{[}\tilde{v}_{k+1}^{*}\Bigl{(}S_{k+1},W_{k+1}\Bigr{)}$
$\Big{|}\tilde{S}_{k},\tilde{\pi}_{k}(\tilde{S}_{k})\biggr{]}\Biggr{\}}\bigg{|}S_{k}\Biggr{]}$
$\stackrel{{\scriptstyle(a)}}{{\geq}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{\tilde{S}_{k+1}}\biggl{[}\tilde{v}_{k+1}^{*}(S_{k+1},W_{k+1})$
$\Big{|}\tilde{S}_{k},\pi_{k}(S_{k})\biggr{]}\bigg{|}S_{k})\Biggr{]}\Biggr{\}}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
$\bigl{(}f^{\rm S}(\tilde{S}_{k},\pi_{k}(S_{k})),W_{k+1}\bigr{)}\Big{|}\tilde{S}_{k},\pi_{k}(S_{k})\Bigr{]}\bigg{|}S_{k}\Biggr{]}\Biggr{\}},$
=$\stackrel{{\scriptstyle(c)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
$\bigl{(}f^{\rm S}(\tilde{S}_{k},\pi_{k}(S_{k})),W_{k+1}\bigr{)}|S_{k+1},\tilde{S}_{k},\pi_{k}(S_{k})\Bigr{]}\Bigr{]}\bigg{|}S_{k}\Biggr{]}\Biggr{\}},$
=$\stackrel{{\scriptstyle(d)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
$\bigl{(}f^{\rm S}(\tilde{S}_{k},\pi_{k}(S_{k})),W_{k+1}\bigr{)}|S_{k+1}\Bigr{]}\Bigr{]}\bigg{|}S_{k}\Biggr{]}\Biggr{\}},$
=$\stackrel{{\scriptstyle(e)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}(S_{k+1})\Bigr{]}$
$\bigg{|}S_{k}\Biggr{]}\Biggr{\}},$
=$\stackrel{{\scriptstyle(f)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}(S_{k+1})\Bigr{]}$
$\bigg{|}S_{k},\pi_{k}(S_{k})\Biggr{]}\Biggr{\}},$
=$\stackrel{{\scriptstyle(g)}}{{=}}\max_{\pi_{k}(S_{k})}\Bigl{\{}\mathrm{E}_{W_{k}}\bigl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\tilde{v}_{k+1}^{*}(S_{k+1})|S_{k},\pi_{k}(S_{k})\bigr{]}\Bigr{\}}$
$\stackrel{{\scriptstyle(h)}}{{\geq}}\max_{\pi_{k}(S_{k})}\Bigl{\{}\mathrm{E}_{W_{k}}\bigl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+v_{k+1}^{*}(S_{k+1})|S_{k},\pi_{k}(S_{k})\bigr{]}\Bigr{\}}$
=$\stackrel{{\scriptstyle(i)}}{{=}}v_{k}^{*}(S_{k}),$ (72)
where (a) follows by interchanging the expectation and maximization (since we generally have $\mathrm{E}[\max\{\cdot\}]\geq\max\{E[\cdot]\}$ according to Jensen’s inequality); (b) follows from $\tilde{S}_{k+1}=\{S_{k+1},W_{k+1}\}$, $S_{k+1}=f^{\rm S}(\tilde{S}_{k},\pi_{k}(S_{k}))$, and the property of conditional probability $\mathrm{E}[g(X)|X]=g(X)$; (c) is due to the property of conditional probability $\mathrm{E}_{X}[\mathrm{E}[Y|X]]=\mathrm{E}[Y]$; (d) follows from Proposition 1, so we have $\mathrm{Pr}(W_{k+1}|S_{k+1},\tilde{S}_{k},\pi_{k}(S_{k}))=\mathrm{Pr}(W_{k+1}|S_{k+1})$; (e) follows from the definition of $\tilde{v}_{k+1}^{*}(S_{k+1})$; (f) holds because the expectation $\mathrm{E}_{W_{k}}[\cdot]$ is calculated for every action $\pi_{k}(S_{k})$ in the action space; (g) follows from $\mathrm{E}_{S_{k+1}}[\cdot|S_{k},\pi_{k}(S_{k})]=\mathrm{E}_{W_{k}}[\cdot|S_{k},\pi_{k}(S_{k})]$ since $S_{k+1}=f^{\rm S}(\tilde{S}_{k},\pi_{k}(S_{k}))$; (h) follows from (69); and (i) follows from the Bellman equation for the original SSDP as given in (-B). Thus (65) is proved for all $k$ and the desired results are shown for the finite horizon problems.
For the infinite horizon problems, the same proof above still holds when $K\rightarrow\infty$. Since the Bellman operator is a contraction mapping, the value functions will converge so we have
$\tilde{v}^{*}(s)\geq v^{*}(s),\forall s\in\mathcal{S},$ (73)
which is similar to (65) for the finite horizon problems. Thus, we have $\tilde{J}^{*}=\mathrm{E}_{\tilde{S}_{0}}[\tilde{v}^{*}(\tilde{S}_{0})]=\mathrm{E}_{S_{0}}[\tilde{v}_{0}^{*}(S_{0})]\geq J^{*}=\mathrm{E}_{S_{0}}[v^{*}(S_{0})]$. This completes the proof of Theorem 2.